# On-chip multi-timescale spatiotemporal optical synchronization

### v2

@@ Abstract
I Abstract
Mode-locking mechanisms are key resources in nonlinear optical phenomena, such as micro-ring solitonic states, and have transformed metrology, precision spectroscopy, and optical communication. However, despite significant efforts, mode-locking has not been demonstrated in the independently tunable multi-timescale regime. Here, we vastly expand the nonlinear mode-locking toolbox into multi-timescale synchronization on a chip. We use topological photonics to engineer a 2D lattice of hundreds of coupled silicon nitride ring resonators capable of hosting nested mode-locked states with a fast ( 1 THz) single-ring and a slow ( 3 GHz) topological super-ring timescales. We demonstrate signatures of multi-timescale mode-locking including quadratic distribution of the pump noise with the two-time azimuthal mode dimensions, as expected by mode-locking theory. Our observations are further corroborated by direct signatures of the near-transform-limit repetition beats and the formation of the temporal pattern on the slow timescale. Moreover, we show that these exotic properties of edge-confined mode-locked states are in sharp contrast to bulk and single-ring counterparts and establish a clear pathway for their identification. Our unprecedented demonstration of mode-locking in topological combs unlocks the implementation of lattice-scale synchronization and independently tunable multi-timescale mode-locking phenomena, also the exploration of the fundamental nonlinearity-topology interplay on a chip.
II Introduction
@@
In the linear regime, a propagating beam of light broadens in space or disperses in time, due to diffraction or chromatic dispersion, respectively. However, nonlinearities may create opposite dispersive or diffractive behavior, for which when they exactly counterbalance allows for mode-locking [1]. The intensity-dependent index of refraction in nonlinear media is an early example that enables self-focusing mechanisms to counteract the diffraction-led divergence, resulting in spatially mode-locked optical solitonic solutions [2, 3]. More recently, pivotal breakthroughs were reported by engineering a delicate balance between the nonlinear Kerr effect and dispersion leading to the discovery of mode-locked temporal dissipative Kerr solitons (DKS) [4, 5] following the same dynamics of their spatial counterpart [6]. DKS have enabled many applications [7] in spectroscopy [8], frequency synthesis [9], ranging [10], and optical clockwork [11, 12].
@@
At its core, the mode-locking of a DKS can be understood as a synchronization phenomenon [13]. Recent years have witnessed significant efforts to develop other DKS synchronization schemes such as remote resonator [14, 15], counter-propagative [16], pulse- [17] or continuous-wave driven [12] DKS, among others. However, all these synchronization schemes rely on the single, or harmonic, frequency grid defined by the DKS repetition rate. The exploration of independently tuneable multi-timescales and multi-spatial modes for which phase-locking may happen has remained scarce [18, 19, 20]. This limitation is partly due to single-resonator systems since they require harmonic synchronization even when dealing with different timescales, such as DKS-breather synchronization [21].
@@
A potential solution is to create nested resonant structures with uncoupled different timescales, allowing synchronization and mode-locking to coexist in a multi-dimensional space. Long 1D chains of microring resonators were theoretically studied [22, 23] as a potential candidate, however, their experimental use is severely limited due to resonant frequency inhomogeneity [24].
@@
Recently, nested-DKS in large 2D arrays, which supports topologically protected edge-mode forming a low-loss super-ring, was proposed [25], followed by the experimental observation of modulation-instability combs in such 2D lattices [26]. Thus, optical synchronization and multi-timescale mode-locked states should theoretically exist but have yet to be experimentally demonstrated.
@@
Here, we overcome these challenges using topological photonics and report the first demonstration of multi-timescale spatiotemporal mode-locking in a 2D array of more than one hundred coupled silicon nitride (SiN) ring resonators. Specifically, using the Kerr effect, we generate optical frequency combs at the edge of the topological lattice and present several direct experimental signatures of their simultaneous mode-locking in the fast ( 1 THz) single-ring and slow ( 3 GHz) topological super-ring timescales. We directly demonstrate the quadratic dependence of the pump noise with the two-time azimuthal mode dimension, as expected by mode-locking theory [27, 28]. Our observations are further substantiated with clear signatures on the slow-time scale with near-transform-limit repetition beats and the formation of temporal patterns. Moreover, we demonstrate that these mode-locking signatures are unique to combs generated in the topological edge bands, and are in sharp contrast to the bulk counterparts, for which we establish a comprehensive pathway for their identification.
III Concept
III.1 The topological lattice
@@
Figure 1 illustrates our system. The topological device consists of a 2D array of coupled ring resonators and is designed to simulate the anomalous quantum Hall (AQH) model for photons [29, 30, 25]. The design, linear dynamics, Hamiltonian description, and band structure of the system are presented in detail in the supplementary information (SI) sections S1-S4. Here we only highlight the key characteristics of our device that are central to understanding the multi-timescale topological mode-locking mechanism.
@@
One of the key characteristics of such a photonic AQH lattice is the existence of an edge band, spectrally located between two bulk bands, near each of the longitudinal mode resonances of the individual rings , as shown in Figure 1b. This repeating bulk-edge-bulk spectrum is numerically calculated by simulating weak broadband light coupled into the device’s input port while monitoring the drop port. Another key feature of the system is that in sharp contrast to the bulk modes, the edge mode resonances within the edge band are spatially confined to the boundary of the lattice with a well-defined round trip time, as shown in Figure 1c. The third important feature is the existence of two timescales in the system, which arise from the fast and slow round trip times associated with individual ring and the super ring, respectively.
III.2 Formation and detection of topological mode-locked states
@@
Figure 1a illustrates our scheme for the generation of nested topologically mode-locked states of light. The starting point is sending a pulsed pump with a 5 ns pulse width and 4 s period (250 kHz repetition rate), in resonance with the edge band of the lattice, into the input port of the device. The Kerr nonlinearity of SiN leads to the formation of frequency combs through a stimulated four-wave mixing process [26]. We note that using a 5 ns pulsed pump provides the peak power necessary for reaching the high comb generation threshold in such a large-scale resonant structure while maintaining a low average power to avoid unwanted thermal effects. Importantly, 5 ns is at least an order of magnitude longer than the slowest time scale in our device, that is the 400 ps, constituting a quasi-continuous wave excitation regime for our experiment.
@@
Similar to single-ring counterparts [5], when the nonlinearity and the dispersion in the system are delicately balanced, the topological optical frequency combs can become mode-locked [25]. In this scenario, the output of the lattice is pulses of light with periodic patterns. Importantly, in contrast to the conventional single-ring microcombs, the periodicity of the output pulses has two timescales, one originating from mode-locking in the fast timescale , and one from the slower timescale . We note that, importantly, these two timescales can be tuned independently (see the SI section S2 for details). Signatures of such two-timescale periodicity can be detected directly by analyzing the output port of a photodiode as shown in Figure 1a.
@@
For detection, the generated frequency combs are analyzed using an OSA, revealing the comb spacing of the single rings and the nested comb spacing of the super ring . The combs are also recorded with a PD and analyzed with an oscilloscope and electrical spectrum analyzer (ESA), respectively, revealing the repetition rate of the combs and their time domain dynamics.
@@
Moreover, in contrast to conventional single rings, our topological lattice gives rise to a band structure with distinct edge and bulk bands. For our AQH lattice, as shown in Figure 1b, this is in the form of an edge band (green) with two bulk bands (gray) on its side, with the width of the bands and the number of edge modes set by the lattice parameters (see the SI sections S1-S3 for details). Importantly, only the edge band has a well-defined round-trip time, and the bulk bands lack such character. As such, temporal mode-locking can only be expected for the edge combs. Throughout this work, we investigate the characteristics of mode-locking in the edge part of the spectrum and show how they significantly differ compared to their bulk counterparts.
IV Optical Signatures of Edge and Bulk frequency combs
@@
IV.1 Optical spectrum profiles
@@
To classify the differences between edge and bulk combs, we focus on a longitudinal mode at 1547.8 nm. We note that coupling the light into the input port selectively only excites the counter-clockwise (CCW) edge modes, which is the edge mode we focus on from here on. We choose the wavelength range of the pump, starting from the bulk at the shorter-wavelength part of the spectrum (1547.6 nm) up to the longer-wavelength part (1548.1 nm), to cover the entire edge band as well as parts of the adjacent two bulk bands. The complete experimental linear characterization of the device is presented in the SI section S4. First, the generated combs are sent to a low-resolution (40 pm), broadband (1300 nm to 1800 nm) OSA. The results are shown in Figure 2a. We observe that in contrast to the bulk which does not have any appreciable comb pattern, pumping the edge band of the lattice generates a comb with a triangular-like envelope. The spacing between the adjacent comb teeth is about 6.3 nm, which is the single-ring free spectral range . The variation of the broadness of the edge combs is related to the linear transmission power at different parts of the edge band, as shown in detail in the SI sections S4 and S13.
IV.2 Spatial Profiles and OPO Threshold Analysis
@@
To show that the combs inherit the topological properties of the linear system and are indeed confined to the boundary of the lattice, we perform direct imaging of the generated bulk and edge combs. While the system is designed to be well-confined in-plane, there is a certain amount of out-of-plane scattering caused by fabrication imperfections and disorder. The light scattered due to surface roughness is collected from above with a 10x objective lens and imaged on an infrared (IR) camera. In addition, we use a 1580 nm long-pass filter to remove the pump and only collect part of the generated comb light.
@@
Figure 2b shows the measured spatial intensity profiles of three types of generated frequency combs. The middle panel shows an edge comb. We observe that the generated comb light is confined to the edge of the lattice and light travels from the input to the output port in the CCW direction. We note that the lattice is strongly coupled to the bus waveguides, therefore the light only weakly continues to circulate around the lattice after reaching the first output port in its path (see the SI section S6 for simulated dynamics in different coupling regimes). Moreover, the propagation is robust and no noticeable scattering into the bulk is observed from the two sharp 90∘ corners. These characteristics show that the comb teeth are indeed generated within the topological edge band and that the topology is preserved even in the presence of strong nonlinearity.
@@
In sharp contrast to the edge comb, when we excite the lattice in the bulk bands, the spatial intensity distribution of generated comb light exhibits no confinement and occupies the bulk of the lattice. The results for two bulk combs are shown in the top and bottom panels of Figure 2b. Additionally, Figure 2b shows simulated spatial distributions of edge and bulk modes in the linear regime for comparison. The details of nonlinear imaging and simulation for many bulk and edge combs are presented in the SI sections S6 and S7.
@@
IV.3 Nestedness of the comb and optical linewidth analysis
@@
To show the nested structure of the edge combs within each comb tooth, we measure the comb output at the drop port using a narrower band (1520nm to 1630 nm) but ultra-high resolution (0.04 pm) heterodyne-based optical spectrum analyzer. The results for a few of the selected comb teeth are shown in Figure 2d. Within each of these comb teeth, we observe the oscillation of another set of well-resolved modes that correspond to the individual edge modes. The spacing between the oscillating edge modes is about 20 pm, which corresponds to the free spectral range of the super-ring formed by the edge states and agrees with linear measurements. Importantly, we observe the nestedness of the comb teeth throughout the edge band. Moreover, we observe that the bulk combs lack this nested structure. See SI sections S9 and S14 for the details of optical linewidth analysis and error estimation.
@@
Next, we perform an optical linewidth analysis of the individual nested comb teeth, which can indicate insights about the noise level of the comb states [27, 31]. We note that our pump has an optical linewidth (2.77 pm with an estimated error of pm, indicated as the red shaded area in Figure 2e, also see the SI sections S9 and S10 for details) that is above the resolution of our high-resolution (0.04 pm) OSA, which enables accurate monitoring of the variation of the optical linewidth of the edge and bulk combs compared to the pump throughout this work. Figure 2e shows the linewidth of the edge modes for the longitudinal mode when the pump wavelength is set near the center of the edge band at 1547.82 nm. We observe around an order of magnitude linewidth reduction for the comb tooth linewidth (blue circles) compared to the cold-cavity counterparts (red circles) which were obtained from the linear drop port transmission measurements. The red dashed line indicates the pump’s optical linewidth. We note that the comb tooth linewidth is close and at times lower than the pump’s linewidth.
@@
Moreover, we perform the optical linewidth analysis for the full pump wavelength range for an individual edge mode. The results for the nested tooth (wavelength 1553.845 nm) are shown in Figure 3f. We observe that throughout the edge band, the nested tooth’s linewidth is close to the pump’s linewidth, and it diverges as the pump is detuned to the bulk on either side. Importantly, this observation indicates that the narrowness of the optical linewidths of the edge combs is predominantly set by the pump’s linewidth. We note that the competition between different noise sources can sometimes lead to comb tooth linewidths becoming even narrower than the pump to a minimum value known as the “fixed point” [27], however, the error values in our data are not sufficiently low to meticulously investigate this effect. Detailed analysis of nesting across the entire wavelength range and over different teeth is presented in the SI section S8.
V Signatures of mode-locking in slow and fast timescales
@@
V.1 Simultaneous mode-locking in fast and slow timescales
Unlike the slow super-ring time scale of 2.5 GHz, direct detection of signatures of the fast time scale of the single-ring round trip time (750 GHz) is beyond the capability of typical oscilloscopes and ESAs. Nevertheless, the elastic tape model [27, 28] provides an accurate theoretical description of optical signatures of mode-locked microcombs that can be measured via optical linewidth analysis. Specifically, this model shows the comb tooth linewidth increases quadratically as a function of mode number separation with respect to a fixed point where noise is minimized to a level near or even lower than that of the pump [27].
@@ V.1 Simultaneous mode-locking in fast and slow timescales
V.2 Mode-locking in the slow timescale
@@
We note two important observations in Figure 3b. First, the linewidth of the beats in the edge band approaches the transform limit, which indicates the existence of mode-locked comb states with varying noise levels throughout most of the edge band. Importantly, linewidths start to diverge as the pump is detuned to the bulk bands on either side of the edge bands. This is in close agreement with the time domain results shown in Figure 4g. Second, the repetition rates within the edge band vary from 2.5 GHz to 4 GHz. This observation, in part, is understood using the group delay measurement of the device in the linear regime. As shown in Figure 3e, the edge part of the spectrum shows an increase of the round trip time between 300-400 ps, when the pump is detuned from the shorter to longer wavelength part of the edge band. This 25 change in the round trip and its increasing trend is the dominant contributor to the observed 35 reduction of the repetition rate seen in Figure 3b. We note that the bulk has no well-defined round trip time which indicates the absence of well-defined repetition rates of the bulk combs.
@@
V.3 Temporal signatures of mode-locking

### v3

@@ Abstract
Mode-locking is foundational to nonlinear optics, enabling advances in metrology, spectroscopy, and communications. However, it remains unexplored in non-harmonic, multi-timescale regimes. Here, we realize on-chip multi-timescale synchronization using topological photonics. We design a 2D lattice of 261 coupled silicon nitride ring resonators that supports nested mode-locked states with fast ( 1 THz) single-ring and slow ( 3 GHz) topological super-ring timescales. We observe clear signatures of multi-timescale mode-locking, including a quadratic distribution of pump noise across both azimuthal mode families, consistent with theory. These findings are supported by near-transform-limited repetition beats and the emergence of periodic temporal patterns on the slow timescale. The edge-confined states show distinct dynamics from bulk and single-ring modes, enabling clear identification. Our results establish topological frequency combs as a robust platform for independently tunable, lattice-scale synchronization, opening new directions for exploring the interplay of nonlinearity and topology in integrated photonics.
Short title
On-chip two-timescale optical synchronization
Teaser
Simultaneous mode locking on two independently designable timescales is observed in a topological photonic lattice.
Introduction
@@
In the linear regime, a propagating beam of light broadens in space and disperses in time, due to diffraction and chromatic dispersion, respectively. However, nonlinearities may create opposite dispersive or diffractive behavior, for which when they exactly counterbalance allows for mode-locking (?). The intensity-dependent index of refraction in nonlinear media is an early example that enables self-focusing mechanisms to counteract the diffraction-led divergence, resulting in spatially mode-locked optical solitonic solutions (?, ?). More recently, pivotal breakthroughs were reported by engineering a delicate balance between the nonlinear Kerr effect and dispersion leading to the discovery of mode-locked temporal dissipative Kerr solitons (DKS) (?, ?) following the same dynamics as their spatial counterpart (?). DKS have enabled many applications (?) in spectroscopy (?), frequency synthesis (?), ranging (?), and optical clockwork (?, ?).
@@
At its core, the mode-locking of a DKS can be understood as a synchronization phenomenon (?). Recent years have witnessed substantial efforts to develop other DKS synchronization schemes such as remote resonator (?, ?), counter-propagative (?), pulse- (?) or continuous-wave driven (?) DKS, among others. However, all these synchronization schemes rely on the single, or harmonic, frequency grid defined by the DKS repetition rate. The exploration of independently designable multi-timescales and multi-spatial modes for which phase-locking may happen has remained scarce (?, ?, ?). This limitation is partly due to single-resonator systems since they require harmonic synchronization even when dealing with different timescales, such as DKS-breather synchronization (?).
@@
A potential solution is to create nested resonant structures with uncoupled different timescales, allowing synchronization and mode-locking to coexist in a multi-dimensional space. Long 1D chains of microring resonators were theoretically studied (?, ?) as a potential candidate, however, their experimental use is severely limited due to resonant frequency inhomogeneity (?).
@@
Recently, nested-DKS in large 2D arrays, which supports topologically protected edge-mode forming a low-loss super-ring, was proposed (?), followed by the experimental observation of modulation-instability combs in such 2D lattices (?, ?). Thus, optical synchronization and multi-timescale mode-locked states should theoretically exist but have yet to be experimentally demonstrated.
@@
Here, we overcome these challenges using topological photonics and report the first demonstration of multi-timescale spatiotemporal mode-locking in a 2D array of more than one hundred coupled silicon nitride (SiN) ring resonators. Specifically, using the Kerr effect, we generate optical frequency combs at the edge of the topological lattice and present several direct experimental signatures of their simultaneous mode-locking in the fast ( 1 THz) single-ring and slow ( 3 GHz) topological super-ring timescales. We directly demonstrate the quadratic dependence of the pump noise in both single-ring and topological super-ring timescales, as expected by mode-locking theory (?, ?). Our observations are further substantiated with clear signatures on the slow-time scale with near-transform-limit repetition beats on RF spectrum and the formation of periodic temporal patterns. Moreover, we demonstrate that these mode-locking signatures are unique to combs generated in the topological edge bands, and they are in sharp contrast to the bulk counterparts, for which we establish a comprehensive pathway for their identification.
Concept
The topological lattice
@@
Figure 1 illustrates our system. The topological device consists of a 2D array of coupled ring resonators and is designed to simulate the anomalous quantum Hall (AQH) model for photons (?, ?, ?). The design, linear dynamics, Hamiltonian description, and band structure of the system are presented in detail in the supplementary information (SI) sections S1-S4. Here we only highlight the key characteristics of our device that are central to understanding the multi-timescale topological mode-locking mechanism.
@@
One of the key characteristics of such a photonic AQH lattice is the existence of an edge band, spectrally located between two bulk bands, near each of the longitudinal mode resonances of the individual rings , as shown in Figure 1B. This repeating bulk-edge-bulk spectrum is numerically calculated by simulating weak broadband light coupled into the device’s input port while monitoring the drop port. Another key feature of the system is that in sharp contrast to the bulk modes, the edge mode resonances within the edge band are spatially confined to the boundary of the lattice with a well-defined round trip time, as shown in Figure 1C. The third important feature is the existence of two timescales in the system, which arise from the fast and slow round trip times associated with the individual ring and the super ring, respectively.
Formation and detection of topological mode-locked states
@@
Figure 1A illustrates our scheme for the generation of nested topologically mode-locked states of light. The starting point is sending a pulsed pump with a 5 ns pulse duration and 4 s period (250 kHz repetition rate), in resonance with the edge band of the lattice, into the input port of the device. The Kerr nonlinearity of SiN leads to the formation of frequency combs through a stimulated four-wave mixing process (?). We note that using a 5 ns pulsed pump provides the peak power necessary for reaching the high comb generation threshold in such a large-scale resonant structure while maintaining a low average power to avoid unwanted thermal effects. Importantly, 5 ns is at least an order of magnitude longer than the slowest time scales in our device, that is the 300 ps, as well as the transient time 400 ps, constituting a quasi-continuous wave excitation regime for our experiment. The discussion of transient time can be found in SI section S19.
@@
Similar to single-ring counterparts (?), when the nonlinearity and the dispersion in the system are delicately balanced, the topological optical frequency combs can become mode-locked (?). In this scenario, the output of the lattice is pulses of light with periodic patterns. We note that in a more general setting, such periodic patterns don’t have to be hyperbolic secant, that is the solution of solitons. Importantly, in contrast to the conventional single-ring microcombs, the periodicity of the output pulses has two timescales, one originating from mode-locking in the fast timescale , and one from the slow timescale . We note that, importantly, these two timescales can be designed independently (see the SI section S2 for details). Signatures of such two-timescale periodicity can be detected directly by analyzing the output port of a photodiode as shown in Figure 1A.
@@
For detection, the generated frequency combs are analyzed using an OSA, revealing the comb spacing of the single rings and the nested comb spacing of the super ring . We index sing-ring modes with and super-ring modes with . The combs are also recorded with a PD and analyzed with an oscilloscope and electrical spectrum analyzer (ESA), respectively, revealing the repetition rate of the combs and their time domain dynamics.
@@
Moreover, in contrast to conventional single rings, our topological lattice gives rise to a band structure with distinct edge and bulk bands. For our AQH lattice, as shown in Figure 1B, this is in the form of an edge band (green) with two bulk bands (gray) on its side, with the width of the bands and the number of edge modes set by the lattice parameters (see the SI sections S1-S3 for details). Importantly, only the edge band has a well-defined round-trip time, and the bulk bands lack such character. As such, temporal mode-locking can only be expected for the edge combs. Throughout this work, we investigate the characteristics of mode-locking in the edge part of the spectrum and show how they substantially differ compared to their bulk counterparts. Details of the complications, limitations and potential future improvements of our work is presented in the SI section S22.
Results
Optical Signatures of Edge and Bulk frequency combs
@@ Results
Optical spectrum profiles
@@
To classify the differences between edge and bulk combs, we focus on a longitudinal mode at 1547.8 nm. We note that coupling the light into the input port selectively only excites the counter-clockwise (CCW) edge modes, which is the edge mode we focus on from here on. We choose the wavelength range of the pump, starting from the bulk at the shorter-wavelength part of the spectrum (1547.6 nm) up to the longer-wavelength part (1548.1 nm), to cover the entire edge band as well as parts of the adjacent two bulk bands. The complete experimental linear characterization of the device is presented in the SI section S4. First, the generated combs are sent to a low-resolution (40 pm), broadband (1300 nm to 1800 nm) OSA. The results are shown in Figure 2A. We observe that in contrast to the bulk which does not have any appreciable envelope pattern, pumping the edge band of the lattice generates a comb with a triangular-like envelope in wavelength. The spacing between the adjacent comb teeth is about 6.3 nm, which is the single-ring free spectral range . One observation is that the bulk comb is broader than edge combs, while the variation of the broadness of the edge combs is related to the linear transmission power at different parts of the edge band, as shown in detail in the SI sections S4 and S13. Apart from broadness, we show that the threshold of bulk combs is higher than that of edge combs, as shown in SI section S17. The transmission of edge combs is also typically higher than bulk, as shown in SI section S13.
Spatial Profiles and OPO Threshold Analysis
@@
To show that the combs inherit the topological properties of the linear system and are indeed confined to the boundary of the lattice, we perform direct imaging of the generated bulk and edge combs. While the system is designed to be well-confined in-plane, there is a certain amount of out-of-plane scattering caused by fabrication imperfections and disorder. The light scattered due to surface roughness is collected from above with a 10x objective lens and imaged on an infrared (IR) camera. In addition, we use a 1600 nm long-pass filter to remove the pump and only collect part of the generated comb light.
@@
Figure 2B shows the measured spatial intensity profiles of two types of generated frequency combs. The first row shows an edge comb. We observe that the generated comb light is confined to the edge of the lattice and light travels from the input to the output port in the CW direction (see the SI section S6 for simulated dynamics in different coupling regimes). Moreover, the propagation is robust and no noticeable scattering into the bulk is observed from the two sharp corners. These characteristics show that the comb teeth are indeed generated within the topological edge band and that the topology is preserved even in the presence of strong nonlinearity. We note that the images shown here are taken from a copy of the main device that is presented throughout the main text. Comparison of the effect of different optical filters is shown in SI section S20.
@@
In sharp contrast to the edge comb, when we excite the lattice in the bulk bands, the spatial intensity distribution of generated comb light exhibits no confinement and occupies the bulk of the lattice. The result for a representative bulk comb is shown in the second row of Figure 2B. Additionally, Figure 2B shows simulated spatial distributions of edge and bulk modes in the linear regime for comparison. The details of nonlinear imaging and simulation for many bulk and edge combs are presented in the SI sections S6 and S7.
@@
Nestedness of the comb and optical linewidth analysis
@@
To show the nested structure of the edge combs within each comb tooth, we measure the comb output at the drop port using a narrower band (1520nm to 1630 nm) but ultra-high resolution (0.04 pm) heterodyne-based optical spectrum analyzer. The results for a few of the selected comb teeth are shown in Figure 2D. Within each of these comb teeth, we observe the oscillation of another set of well-resolved modes that correspond to the individual edge modes. The spacing between the oscillating edge modes is about 20 pm, which corresponds to the free spectral range of the super-ring formed by the edge states and agrees with linear measurements. See the detailed demonstration of 20pm spacing in (?) as well as SI section S21. Importantly, we observe the nestedness of the comb teeth throughout the edge band. Moreover, we observe that the bulk combs lack this nested structure. See SI sections S9 and S14 for the details of optical linewidth analysis and error estimation.
@@
Next, we perform an optical linewidth analysis of the individual nested comb teeth, which provides insights on the noise level of the comb states (?, ?). We note that our pump has an optical linewidth (2.77 pm with an estimated error of pm, indicated as the red shaded area in Figure 2E, also see the SI sections S9 and S10 for details) that is above the resolution of our high-resolution (0.04 pm) OSA, which enables accurate monitoring of the variation of the optical linewidth of the edge and bulk combs compared to the pump throughout this work. Figure 2E shows the linewidth of the edge modes for the longitudinal mode when the pump wavelength is set near the center of the edge band at 1547.82 nm. We observe around an order of magnitude linewidth reduction for the comb tooth linewidth (blue circles) compared to the cold-cavity counterparts (red circles) which were obtained from the linear drop port transmission measurements. The red dashed line indicates the pump’s optical linewidth. We note that the comb tooth linewidth is close and at times lower than the pump’s linewidth.
@@
Moreover, we perform the optical linewidth analysis for the full pump wavelength range for an individual edge mode. The results for the nested tooth (wavelength 1553.845 nm) are shown in Figure 3F. We observe that throughout the edge band, the nested tooth’s linewidth is close to the pump’s linewidth, and it diverges as the pump is detuned to the bulk on either side. Importantly, this observation indicates that the narrowness of the optical linewidths of the edge combs is predominantly set by the pump’s linewidth. We note that the competition between different noise sources can sometimes lead to comb tooth linewidths becoming even narrower than the pump to a minimum value known as the “fixed point” (?), however, the error values in our data are not sufficiently low to meticulously investigate this effect. Detailed analysis of nesting across the entire wavelength range and over different teeth is presented in the SI section S8.
Signatures of mode-locking in slow and fast timescales
@@
Simultaneous mode-locking in fast and slow timescales
Unlike the slow super-ring time scale of 2.5 GHz, direct detection of signatures of the fast time scale of the single-ring round trip time (750 GHz) is beyond the capability of typical oscilloscopes and ESAs. Nevertheless, the elastic tape model (?, ?) provides an accurate theoretical description of optical signatures of mode-locked microcombs that can be measured via optical linewidth analysis. Specifically, this model which is an ab-initio derivation assuming a fixed repetition rate for a train of pulses, shows the comb tooth linewidth increases quadratically as a function of mode number separation with respect to a fixed point (?).
@@
Mode-locking in the slow timescale
@@
We note two important observations in Figure 3B. First, the linewidth of the ESA beat notes within the edge band approaches the transform limit, which highly suggests the existence of mode-locked comb states. Moreover, this near-transform-limited linewidth persists over a relatively wide portion of the edge band, suggesting a broad range of possible mode-locking states. Importantly, linewidths start to diverge as the pump is detuned to the bulk bands on either side of the edge bands. This is in close agreement with the time domain results shown in Figure 4G. Second, the repetition rates within the edge band vary from 2.5 GHz to 4 GHz. This observation, in part, is understood using the group delay measurement of the device in the linear regime. As shown in Figure 3E, the edge part of the spectrum shows an increase of the round trip time between 300-400 ps, when the pump is detuned from the shorter to longer wavelength part of the edge band. This 25 change in the round trip and its increasing trend is the dominant contributor to the observed 35 reduction of the repetition rate seen in Figure 3B. We note that the bulk has no well-defined round trip time which indicates the absence of well-defined repetition rates of the bulk combs.
@@
Temporal signatures of mode-locking